Find the variance of a discrete random variable X - the number of occurrences of event A in four independent tests, if the probability of occurrence of event A in each test is 0.1. Write in the form of a table the law of distribution of the random variable X


1
Expert's answer
2021-12-27T17:38:02-0500

Solution:

"p=0.1,q=0.9,n=4"

"X\sim Bin(n,p)
\\ Var(X)=npq=4(0.1)(0.9)=0.36"

"P(X=0)=\ ^4C_0(0.1)^0(0.9)^4=0.6561
\\ P(X=1)=\ ^4C_1(0.1)^1(0.9)^3=0.2916
\\ P(X=2)=\ ^4C_2(0.1)^2(0.9)^2=0.0486
\\ P(X=3)=\ ^4C_3(0.1)^3(0.9)^1=0.0036
\\ P(X=4)=\ ^4C_4(0.1)^4(0.9)^0=0.0001"